How would I solve this problem step for step....k/5+4=4/5
Answer by Cromlix(4381)


k/5 + 4 = 4/5
Common denominator = 5
k + 20 = 4
k = 4 - 20
k = -16
Proof:
k/5 + 4 = 4/5
-16/5 + 20/5 = 4/5
4/5 = 4/5
Hope this helps.
;-)